OpenStudy (logan13):

Which is the correct classification for the triangle? A. acute triangle B. right triangle C. obtuse triangle

OpenStudy (logan13):

@KyanTheDoodle

OpenStudy (logan13):

@DanJS

OpenStudy (logan13):

@study_buddy99

OpenStudy (danjs):

IF the sum of the 3 angles in a triangle total to 180 degrees, what is that missing angle?

OpenStudy (logan13):

B

OpenStudy (danjs):

there are 3 interior angles in a triangle, if you add them up, you get 180 degrees, what is that missing angle A A + 35 + 45 = 180

OpenStudy (logan13):

so 35 + 45 - 180

OpenStudy (danjs):

that woul dbe negative A... A + 80 = 180 A = 180 - 80 A=100

OpenStudy (danjs):

Right means the triangle has a corner 90 degree angle, this does not Acute means, no angle is above 90 obtuse means, an angle is above 90

OpenStudy (logan13):

soo c

OpenStudy (danjs):

is that your final answer

OpenStudy (logan13):

yes

OpenStudy (danjs):

right, i hope ya learned sometihng